An object has a momentum of 32.0 kg m/sec in the positive X direction. A force in the negative X direction is applied for 1.3 seconds and the object stops.
a) What is the magnitude of the Force?
b) If the mass of the object is 4.00 kg, what is the initial kinetic energy?
c) What is the Impulse applied to the object to cause it to stop?
d) What is the final momentum of the object?
e) What is the final kinetic energy of the object?
1
Expert's answer
2020-11-16T07:45:25-0500

a) "F=\frac{p}{\triangle t}=\frac{32}{ 1.3}=24.6" N

b) "E=\frac{mv^2}{2}=\frac{(mv)^2}{2m}=\frac{p^2}{2m}=\frac{32^2}{2\cdot4}=128" J

c) "p=32" "\frac{kg\cdot m}{s}"

d) "v=0" "\frac{ m}{s}", so "p=0" "\frac{kg\cdot m}{s}"

e) "v=0" "\frac{ m}{s}", so "E=0" J